# Density-valued time series: Nonparametric density-on-density regression

### v1

@@ Abstract
This paper is concerned with forecasting probability density functions. Density functions are nonnegative and have a constrained integral; they thus do not constitute a vector space. Implementing unconstrained functional time-series forecasting methods is problematic for such nonlinear and constrained data. A novel forecasting method is developed based on a nonparametric function-on-function regression, where both the response and the predictor are probability density functions. Through a series of Monte-Carlo simulation studies, we evaluate the finite-sample performance of our nonparametric regression estimator. Using French departmental COVID19 data and age-specific period life tables in the United States, we assess and compare finite-sample forecast accuracy between the proposed and several existing methods.
Keywords: Bayes space; Convex constraint; Density forecasting; Nonparametric function-on-function regression; Time series of densities.
@@ 1 Introduction
Recent advances in computer storage and data collection have enabled researchers to record data of characteristics varying over a continuum (time, space, depth, wavelength, etc.). In diverse branches of science, the data are collected by a spectrometer, rain gauges, electroencephalographs, or even a high-performance computer. In all these cases, a number of subjects are observed densely over time, space, or both. Through the application of interpolation or smoothing techniques, these data become functions that can be represented as a curve, image, or shape. Functional data analysis has arisen as a field of statistics that provides statistical tools for analyzing this type of information. For an overview of functional data analysis, Ramsay & Silverman (2005) presented several state-of-the-art statistical techniques while Ferraty & Vieu (2006) listed a range of nonparametric techniques. Hsing & Eubank (2015) provided theoretical foundations with an introduction to linear operators, while Bosq (2000) studied functional autoregressive process from a linear operator perspective. An introduction to temporally dependent functional data can be found in Kokoszka & Reimherr (2017), while Mateu & Giraldo (2022) studied spatially-dependent functional data.
@@
Functional data that are samples of one-dimensional random probability density functions (PDFs) are common. Examples include income distributions (Kneip & Utikal, 2001), distributions of the times when bids are submitted in an online auction (Jank et al., 2008), functional connectivity in the brain (Petersen & Müller, 2019), distribution of image features from head CT scans (Salazar et al., 2020), distributions of stock return (Harvey et al., 2016; Bekierman & Gribisch, 2021; Zhang et al., 2022), the age distribution of fertility rates (Mazzuco & Scarpa, 2015) and age distribution of life-table death counts in demography (Bergeron-Boucher et al., 2017).
@@
To model density-valued functional data, Jones & Rice (1992) estimated density function by kernel density estimator and displayed density function via principal components. Nerini & Ghattas (2007) proposed regression trees where the response variable is PDFs. van der Linde (2008) introduced Bayesian functional principal component analysis and applied it to simulated data consisting of nonparametric density estimates. Srivastava et al. (2011) proposed a time-warping function where the square root transformation of densities resides in the Hilbert space. Dai (2022) considered statistical inference of the Fréchet mean on the Hilbert sphere with application to random densities. Zhang et al. (2022) proposed Wasserstein autoregressive models for density time series. Shang & Haberman (2020, 2025a, 2025b) considered the center log-ratio transformation,
@@
While any PDF $f$ may be thought of as random element of a Hilbert space, it does not reside in linear space due to summability and non-negativity constraints (i.e. $f\geq 0$ and $\int f=1$); if $f$, $g$ are two PDFs and $a$, $b$ two real values, then the combination $a\,f\,+\,b\,g$ is not a PDF anymore. Consequently, the standard space of square-integrable functions $\mathcal{L}^{2}$ is not the most appropriate for representing PDFs. A natural way to deal with such constraints is to peel them away by an invertible transformation that maps densities onto a linear space. From an extrinsic viewpoint, such transformations include the Bayes Hilbert space approach (see, e.g., Hron et al., 2016; Menafoglio et al., 2022), compositional data analysis (CoDa) (Shang & Haberman, 2020), cumulative distribution function transformation (Shang & Haberman, 2025a), $\alpha$ transformation (Shang & Haberman, 2025b) and log quantile density transformation (LQDT) (see, e.g., Petersen & Müller, 2016; Han et al., 2020). Kokoszka et al. (2019) compared the finite-sample performance of density estimation and forecasting among the LQDT, CoDa, unconstrained functional principal component regression of Horta & Ziegelmann (2018), and a skewed-$t$ distribution of Wang (2012). Kokoszka et al. (2019) recommended the LQDT and CoDa, two benchmark methods in our empirical studies.
@@
In this paper, we focus on a random process taking values in some space of densities. We aim to forecast nonparametrically a future density given those observed in the past. Density forecasting is of great importance, especially in demography. In the realm of actuarial statistics, an important use of density forecasting is to predict the age distribution of life-table death counts as a means of computing survival probabilities. In such a model, any wrong specification of the forecast density may produce an inaccurate estimate of value-at-risk and make the asset holder, i.e., insurance companies, unable to control risk (see, e.g., Shang, 2013). In Figure 1, we present a time series of age-specific life-table death counts for the United States of America (USA) from 1933 to 2023 obtained from the Human Mortality Database (2025).
@@
This setting faces three non-standard frameworks: 1) nonparametric density-on-density regression, 2) the dependency of the sequence of observed density functions, and 3) the convexity property of the space of densities that do not form a linear space. To take into account the nonlinear feature of the space of PDFs, we consider the Bayes Hilbert space approach (see, e.g., Egozcue et al., 2006; Van den Boogaart et al., 2010, 2014; Hron et al., 2016). The Bayes Hilbert spaces provide a mathematical environment allowing the extension of the Aitchison geometry from finite-dimensional compositional data (Aitchison, 1986; Pawlowsky-Glahn et al., 2015) to an infinite-dimensional setting. The Bayes Hilbert space is a linear space of PDFs designed for statistical models involving random PDFs (see, e.g., Delicado, 2011; Hron et al., 2016; Seo & Beare, 2019; Joen & Park, 2020).
@@
A nonparametric density-on-density regression, where the response is one step ahead of the predictor in a time series of densities, is defined in the Bayes Hilbert space. A distance-based approach with a kernel-type estimator is proposed. Its flexibility makes the implementation attractive as various distance metrics may be considered. Its simplicity also makes the implementation easy and computationally fast as it does in the more standard nonparametric function-on-function regression setting (see, e.g., Ferraty et al., 2012; Lian, 2012). The advantage of nonparametric modelling is to capture the possible nonlinear dependence between two consecutive random PDFs. Unfortunately, instead of observing the density-valued random process, we sometimes only have samples drawn from these density functions (see also Kokoszka et al., 2019; Shang & Haberman, 2020; Zhang et al., 2022). This is why our methodology often includes, in its first step, the estimation of the densities before estimating the nonparametric relationship between densities.
@@
The paper is organised as follows. Section 2.1 reminds the basic geometrical concepts of Bayes Hilbert space. In Section  2.2, we describe a model, estimation and forecasting scheme for our nonparametric density-on-density regression. In Section 2.3, we present asymptotic properties associated with our Bayes Nadaraya-Watson (NW) estimator. In Section 3.1, we describe how we estimate density via a kernel density estimator, the selection of optimal bandwidth, and the criterion used to assess density estimation accuracy. In Section 3.2, we evaluate and compare our Bayes NW estimator in the nonparametric density-on-density regression via a series of simulation studies. A French COVID19 hospitalisation data analysis is shown in Section 3.3, and a data analysis of age-specific life-table death count in the United States (U.S.) is presented in Section 3.4. The conclusion is given in Section 4, along with some ideas on how the methodology can be further extended.
@@ 2.1 Bayes Hilbert space: a vector space of PDFs
Let $\mathcal{B}^{2}(I)$ be the set of bounded PDFs $f$ with continuous support $I=[a,\,b]$ for which the natural logarithm is square integrable:
$\mathcal{B}^{2}(I)=\left\{f:I\mapsto]0,\,+\infty[,\,\int_{I}f=1,\,\int_{I}\{\ln(f)\}^{2}\,<\,+\infty\right\}.$
Remark: to simplify notations, when there is no ambiguity, arguments of integrand are forgotten. Two fundamental operators $\oplus$ and $\odot$ respectively called perturbation and powering are defined for any $f$, $g$ in $\mathcal{B}^{2}(I)$ and $r$ in $\mathbb{R}$:
[…]
@@
Let ${\bf 0}_{\mathcal{B}}=\frac{1}{b-a}1_{I}$, the uniform PDF defined on $I$, is the neutral element ($f\oplus{\bf 0}_{\mathcal{B}}={\bf 0}_{\mathcal{B}}\oplus f=f$) for perturbation and $f^{-1}$ is the unique element such that $f\oplus f^{-1}=f^{-1}\oplus f={\bf 0}_{\mathcal{B}}$. The power is associative ($r\odot(s\odot f)=(r\,s)\odot f$), distributive with respect to perturbation ($r\odot(f\oplus g)=(r\odot f)\oplus(r\odot g))$ and scalar addition ($(r+s)\odot f=(r\odot f)\oplus(s\odot f)$), and 1 is its neutral element. Based on the perturbation and powering, a third operator $\ominus$ called perturbation-substraction is defined:
@@
In $\mathcal{B}^{2}(I)$, the inner product is defined as:
@@
and one has useful properties: 1) $(\mathcal{B}^{2}(I),\,\oplus,\,\odot,\,\langle\cdot,\cdot\rangle_{\mathcal{B}})$is a vector space, 2) $(\mathcal{B}^{2}(I),\,\oplus,\,\odot,\,\langle\cdot,\cdot\rangle_{\mathcal{B}})$forms a separable Hilbert space called Bayes Hilbert space (see Egozcue et al., 2006; Van den Boogaart et al., 2014). At last, for any $f$, $g$ in $\mathcal{B}^{2}(I)$, $\|f\|_{\mathcal{B}}=\sqrt{\langle f,f\rangle_{\mathcal{B}}}$ is the norm associated with the inner product. One can remark that $\|f\ominus g\|_{\mathcal{B}}=0$ is equivalent to $f\ominus g={\bf 0}_{\mathcal{B}}$, which results in $f(u)=g(u)$ for any $u\in I$.
Isometry between $\mathcal{B}^{2}(I)$ and $\mathcal{L}^{2}_{0}(I)$.
Let $\mathcal{L}^{2}_{0}(I)$ be the space of square-integrable real functions $f$ on $I$ with $\int_{I}f=0$. There exists an isometric isomorphism between $\mathcal{B}^{2}(I)$ and $\mathcal{L}^{2}_{0}(I)$ called centered $\log$-ratio (clr) transformation, defined for any PDF $f$ in $\mathcal{B}^{2}(I)$ and any $u$ in $I$:
@@
Once the Bayes Hilbert space framework is reminded, one can introduce our nonparametric density-on-density regression when considering a time series of PDFs.
@@ 2.2.1 Density-on-density regression model
Let $\{\mathscr{F}_{t}\}_{t\in\mathbb{Z}}$ be a stochastic process taking values in $\mathcal{B}^{2}(I)$ and consider $N+1$ random PDFs $(f_{1},\ldots,f_{N+1})$ as realizations of $(\mathscr{F}_{1},\ldots,\mathscr{F}_{N+1})$. The main goal is to forecast a future PDF $f_{N+2}$ given those observed in the past ($f_{N+1},\,f_{N},\ldots,f_{1}$). To this end, we propose the nonparametric density-on-density regression model
@@
where $m(\cdot)$ is an unknown smooth operator mapping $\mathcal{B}^{2}(I)$ into $\mathcal{B}^{2}(I)$, and for each $t$, $\text{E}(\epsilon_{t}|f_{t})={\bf 0}_{\mathcal{B}}$, where the conditional expectation is defined for $\mathcal{B}^{2}(I)$-valued random PDFs. The definition of conditional expectation for Hilbert-valued random variable can be found in Bosq (2000). Under Model (1), it is easy to check that, for each $t$, $\text{E}(\epsilon_{t}|f_{t})={\bf 0}_{\mathcal{B}}$ is equivalent to $m(f)=\text{E}\left(f_{t+1}|f_{t}=f\right)$ (see Section A in Appendix). Our main challenge is to define for any given density $f$ an estimator $\widetilde{m}(f)$ of the unknown regression operator $m(f)=\text{E}\left(f_{t+1}|f_{t}=f\right)$ fulfilling density features (nonnegative functions that integrate to one) and good asymptotic behaviour while no reducing functional form (i.e. linearity, additivity, functional index model, etc) of the unknown regression operator $m(\cdot)$ is assumed.
@@ 2.2.2 Kernel estimator
With some smoothness properties, the functional form of $m(\cdot)$ is often estimated in a data-driven manner. There are a number of nonparametric functional estimators, such as functional Nadaraya-Watson (NW) estimator (Ferraty & Vieu, 2006), functional local linear estimator (Berlinet et al., 2011; Ferraty & Nagy, 2022), functional $k$-nearest neighbour estimator (Burba et al., 2009, for scalar response and Lian, 2011, for functional response), functional smoothing splines (Crainiceanu & Goldsmith, 2010), functional wavelet estimator (Antoniadis & Sapatinas, 2003), distance-based local linear estimator (Boj et al., 2010) and functional neural network (Rossi & Villa, 2006). Throughout the paper, we introduce the functional Bayes NW estimator, i.e., the functional NW estimator in the Bayes-Hilbert space, because of its simplicity and mathematical elegance. The essence of functional Bayes NW kernel smoothing is to allow flexible estimation of the unknown regression operator. The functional Bayes NW estimator of the conditional mean can be defined as
@@
$\widetilde{m}_{N}(f)\,=\,\bigoplus^{N}_{t=1}w_{h_{r}}(f_{t},\,f)\,\odot\,f_{t+1}\ \mbox{ with }\ w_{h_{r}}(f_{t},\,f)\,=\,\frac{K_{r}(h_{r}^{-1}\,\|f_{t}\ominus f\|_{\mathcal{B}})}{\sum^{N}_{t=1}K_{r}(h_{r}^{-1}\,\|f_{t}\ominus f\|_{\mathcal{B}})},$ (2)
@@
where $K_{r}(\cdot)$ is an asymmetric kernel function and $h_{r}$ is a positive smoothing parameter (i.e. bandwidth). The vector space property of $\mathcal{B}^{2}(I)$ ensures that the Bayes NW estimator is a PDF. Depending on the kernel function and bandwidth parameter, the nonparametric estimator in (2) is a weighted average of the observed densities and has two interesting properties: it fulfils the PDF constraints by construction, and due to its simplicity, its implementation is straightforward and computationally fast.
@@
In practice, conditionally on the densities $f_{1},\ldots,f_{N+1}$, we observe $n_{t}$ copies $X_{t,1},\ldots,X_{t,n_{t}}$ that share the same marginal density $f_{t}$ for each $t=1,\ldots,N+1$. In other words, the PDFs are not directly observable. This situation is particularly common in epidemiology. For instance, suppose that at each date $t$, COVID-19 confirmed cases corresponding to $\omega$ French departments are observed. In this case, $X_{t,j}$ is the COVID-19 confirmed cases at date $t$ of the department $j$ and all the sample sizes are equal $(n_{1}=n_{2}=\dots=\omega)$. This is why our methodology first focuses on kernel estimators $\widehat{f}_{1},\ldots,\widehat{f}_{N+1}$ of the $N+1$ probability densities $f_{1},\ldots,f_{N+1}$. In the second step, we just plug the estimated densities $\widehat{f}_{1},\ldots,\widehat{f}_{N+1}$ into the nonparametric density-on-density regression estimator (2) which provides our definitive kernel estimation:
$\widehat{m}_{N}(f)\,=\,\bigoplus^{N}_{t=1}w_{h_{r}}(\widehat{f}_{t},\,f)\,\odot\,\widehat{f}_{t+1},$
@@
where $w_{h}(\widehat{f}_{t},\,f)$ is defined in (2). A standard by-product of this estimating procedure in the time series context is the following one-step-ahead sequential forecasting scheme. From the estimated density functions $\widehat{f}_{1},\ldots,\widehat{f}_{N+1}$, we compute $\widehat{f}_{N+2}=\widehat{m}_{N}(\widehat{f}_{N+1})$ which is the prediction of the density $f_{N+2}$. To forecast the next density, just consider the sample of random densities completed with the predicted one (i.e. $\widehat{f}_{1},\ldots,\widehat{f}_{N+1},\widehat{f}_{N+2}$) and build $\widehat{f}_{N+3}=\widehat{m}_{N+1}(\widehat{f}_{N+2})$. This one-step-ahead forecasting scheme can be iterated to get $\widehat{f}_{N+\mathcal{H}}$ at any reasonable forecast horizon $\mathcal{H}$.
2.2.3 Dependence model

### v2

@@ Abstract
This paper is concerned with forecasting probability density functions. Density functions are nonnegative and have a constrained integral; thus, they do not constitute a vector space. Implementing unconstrained functional time-series forecasting methods is problematic for such nonlinear and constrained data. A novel forecasting method is developed based on a nonparametric function-on-function regression, where both the response and the predictor are probability density functions. Asymptotic properties of our nonparametric regression estimator are established, as well as its finite-sample performance through a series of Monte-Carlo simulation studies. Using COVID-19 data from the French department and age-specific period life tables from the United States, we assess and compare the finite-sample forecast accuracy of the proposed method with several existing methods.
Keywords: Bayes space; Convex constraint; Density forecasting; Kernel estimator; MISE; Nonparametric function-on-function regression; Time series of densities.
@@ 1 Introduction
Recent advances in computer storage and data collection have enabled researchers to record data of characteristics varying over a continuum (time, space, depth, wavelength, etc.). In various branches of science, data are collected using a spectrometer, rain gauges, electroencephalographs, or even a high-performance computer. In all of these cases, a number of subjects are observed densely over time, space, or both. Through the application of interpolation or smoothing techniques, these data become functions that can be represented as a curve, image, or shape. Functional data analysis has arisen as a field of statistics that provides statistical tools for analysing this type of information. For an overview of functional data analysis, Ramsay & Silverman (2005) presented several state-of-the-art statistical techniques, while Ferraty & Vieu (2006) listed a range of nonparametric techniques. Hsing & Eubank (2015) provided theoretical foundations with an introduction to linear operators, while Bosq (2000) studied the functional autoregressive process from a linear operator perspective. An introduction to temporally dependent functional data can be found in Kokoszka & Reimherr (2017), while Mateu & Giraldo (2022) studied spatially-dependent functional data.
@@
Functional data that are samples of one-dimensional random probability density functions (PDFs) are common. Examples include income distributions (Kneip & Utikal, 2001), distributions of times when bids are submitted in an online auction (Jank et al., 2008), functional connectivity in the brain (Petersen & Müller, 2019), distribution of image features from head CT scans (Salazar et al., 2020), distributions of stock returns (Harvey et al., 2016; Bekierman & Gribisch, 2021; Zhang et al., 2022), the age distribution of fertility rates (Mazzuco & Scarpa, 2015), and age distribution of life-table death counts in demography (Bergeron-Boucher et al., 2017).
@@
To model density-valued functional data, Jones & Rice (1992) estimated the density function by a kernel density estimator and displayed the density function via principal components. Nerini & Ghattas (2007) proposed regression trees where the response variable is PDF. van der Linde (2008) introduced Bayesian functional principal component analysis and applied it to simulated data consisting of nonparametric density estimates. Srivastava et al. (2011) proposed a time-warping function where the square root transformation of densities resides in Hilbert space. Dai (2022) considered statistical inference of the Fréchet mean in the Hilbert sphere with application to random densities. Zhang et al. (2022) proposed Wasserstein autoregressive models for density time series. Shang & Haberman (2020, 2025a, 2025b) considered the centred log-ratio transformation,
@@
While any PDF $f$ may be thought of as a random element of a Hilbert space, it does not reside in linear space due to summability and non-negativity constraints (i.e. $f\geq 0$ and $\int f=1$); if $f$, $g$ are two PDFs and $a$, $b$ two real values, then the combination af+bga\,f\,+\,b\,g is no longer a PDF (except in the special case b=1ab=1-a with $a\geq 0$). Consequently, the standard space of square-integrable functions $\mathcal{L}^{2}$ is not the most appropriate for representing PDFs. A natural way to deal with such constraints is to peel them away by an invertible transformation that maps densities onto a linear space. From an extrinsic viewpoint, such transformations include the Bayes Hilbert space approach (see, e.g., Hron et al., 2016; Menafoglio et al., 2022), compositional data analysis (CoDa) (Shang & Haberman, 2020), cumulative distribution function transformation (Shang & Haberman, 2025a), $\alpha$ transformation (Shang & Haberman, 2025b) and log quantile density transformation (LQDT) (see, e.g., Petersen & Müller, 2016; Han et al., 2020). Kokoszka et al. (2019) compared the finite-sample performance of density estimation and forecasting among the LQDT, CoDa, unconstrained functional principal component regression of Horta & Ziegelmann (2018), and a skewed-$t$ distribution of Wang (2012). Kokoszka et al. (2019) recommended the LQDT and CoDa, two benchmark methods in our empirical studies.
@@
In this paper, we focus on a random process that takes values in a space of densities. We aim to forecast the future density nonparametrically, given those observed in the past. Density forecasting is of great importance, especially in demography. In the realm of actuarial statistics, an important use of density forecasting is to predict the age distribution of life-table death counts as a means of computing survival probabilities. In such a model, any incorrect specification of the forecast density can produce an inaccurate estimate of the value-at-risk and make the asset holder, i.e., insurance companies, unable to control risk (see, e.g., Shang, 2013). In Figure 1, we present a time series of age-specific life-table death counts for the United States of America (USA) from 1933 to 2023 obtained from the Human Mortality Database (2025).
@@
This setting faces three non-standard frameworks: 1) nonparametric density-on-density regression, 2) the dependence of the sequence of observed density functions, and 3) the convexity property of the space of densities that do not form a linear space. To take into account the nonlinear feature of the space of PDFs, we consider the Bayes Hilbert space approach (see, e.g., Egozcue et al., 2006; Van den Boogaart et al., 2010, 2014; Hron et al., 2016). The Bayes Hilbert spaces provide a mathematical environment that allows the extension of Aitchison geometry from finite-dimensional compositional data (Aitchison, 1986; Pawlowsky-Glahn et al., 2015) to an infinite-dimensional setting. The Bayes Hilbert space is a linear space of PDFs designed for statistical models involving random PDFs (see, e.g., Delicado, 2011; Hron et al., 2016; Seo & Beare, 2019; Joen & Park, 2020)).
@@
A nonparametric density-on-density regression, where the response is one-step-ahead of the predictor in a time series of densities, is defined in the Bayes Hilbert space. A distance-based approach with a kernel-type estimator is proposed. Its flexibility makes the implementation attractive, since various distance metrics may be considered. Its simplicity also makes the implementation easy and computationally fast as it does in the more standard nonparametric function-on-function regression setting (see, e.g., Ferraty et al., 2012a; Lian, 2012). The advantage of nonparametric modelling is to capture the possible nonlinear dependence between two consecutive random PDFs. Unfortunately, instead of observing the density-valued random process, we sometimes only have samples drawn from these density functions (see also Kokoszka et al., 2019; Shang & Haberman, 2020; Zhang et al., 2022). This is why our methodology often includes, in its first step, the estimation of the densities before nonparametrically estimating the relationship between densities.
@@
The paper is organised as follows. Section 2.1 reminds of the basic geometric concepts of the Bayes Hilbert space. In Section 2.2, we describe a model, estimation and forecasting scheme for our nonparametric density-on-density regression. In Section 2.3, we present asymptotic properties associated with our Bayes Nadaraya-Watson (NW) estimator. In particular, we establish a condition that ensures the existence of a stationary PDF-valued random process consistent with the proposed nonparametric PDF-on-PDF regression model. We also establish the asymptotic behaviour of the mean integrated squared error of the kernel density estimator when the target is a PDF-valued random variable. In Section 3.1, we describe how we implement the kernel density estimator, the selection of optimal bandwidth, and the criterion used to assess density estimation accuracy. In Section 3.2, we evaluate and compare our Bayes NW estimator in the nonparametric density-on-density regression via a series of simulation studies. A French COVID-19 hospitalisation data analysis is shown in Section 3.3, and a data analysis of age-specific life-table death count in the United States (U.S.) is presented in Section 3.4. The conclusion is given in Section 4, along with some ideas on how the methodology can be further extended.
@@ 2.1 Bayes Hilbert space: a vector space of PDFs
Let $\mathcal{B}_{2}(I)$ be the set of bounded PDFs $f$ with continuous support I=[a,b]I=[a,\,b] for which the natural logarithm is square integrable:
$\mathcal{B}_{2}(I)=\left\{f:I\mapsto]c,\,+\infty[\mbox{ with }c>0,\,\int_{I}f=1,\,\int_{I}\{\ln(f)\}^{2}\,<\,+\infty\right\}.$
Remark: to simplify notation, we omit the arguments of the integrand whenever no ambiguity arises. Two fundamental operators $\oplus$ and $\odot$ respectively called perturbation and powering are defined for any $f$, $g$ in $\mathcal{B}_{2}(I)$ and $r$ in $\mathbb{R}$:
$(f\oplus g)(t)=\frac{f(t)\,g(t)}{\int_{I}f(u)\,g(u)\,du},\qquad\displaystyle(r\odot f)(t)=\frac{f^{r}(t)}{\int_{I}f^{r}(u)\,du}.$
@@
Let ${\bf 0}_{\mathcal{B}}=\frac{1}{b-a}1_{I}$, the uniform PDF defined on $I$, be the neutral element ($f\oplus{\bf 0}_{\mathcal{B}}={\bf 0}_{\mathcal{B}}\oplus f=f$) for the perturbation and $f^{-1}$ be the unique element such that $f\oplus f^{-1}=f^{-1}\oplus f={\bf 0}_{\mathcal{B}}$. The power is associative ($r\odot(s\odot f)=(r\,s)\odot f$), distributive with respect to the perturbation ($r\odot(f\oplus g)=(r\odot f)\oplus(r\odot g))$ and the scalar addition ($(r+s)\odot f=(r\odot f)\oplus(s\odot f)$), and 1 is its neutral element. Based on perturbation and powering, a third operator $\ominus$ called perturbation-substraction is defined:
@@
In $\mathcal{B}_{2}(I)$, the inner product is defined as:
@@
[…]
Isometry between $\mathcal{B}_{2}(I)$ and $\mathcal{L}^{2}_{0}(I)$.
Let $\mathcal{L}^{2}_{0}(I)$ be the space of square-integrable real functions $f$ on $I$ with $\int_{I}f=0$. There exists an isometric isomorphism between $\mathcal{B}_{2}(I)$ and $\mathcal{L}^{2}_{0}(I)$ called the centred $\log$-ratio (clr) transformation, defined for any PDF $f$ in $\mathcal{B}_{2}(I)$ and any $u$ in $I$:
@@
After recalling the Bayes Hilbert space framework, we introduce our nonparametric density-on-density regression for the analysis of time series of PDFs.
@@ 2.2.1 Density-on-density regression model
Let $\{f_{t}\}_{t\in\mathbb{Z}}$ be a stochastic process taking values in $\mathcal{B}_{2}(I)$ and consider the $N+1$ random PDFs $(f_{1},\ldots,f_{N+1})$. The main goal is to forecast a future PDF $f_{N+2}$ given those observed in the past ($f_{N+1},\,f_{N},\ldots,f_{1}$). To this end, we propose the nonparametric density-on-density regression model
@@
where $m(\cdot)$ is an unknown smooth operator mapping $\mathcal{B}_{2}(I)$ into $\mathcal{B}_{2}(I)$, and for each $t$, $\text{E}(\epsilon_{t}|f_{t})={\bf 0}_{\mathcal{B}}$, where the conditional expectation is defined for $\mathcal{B}_{2}(I)$-valued random PDFs. The definition of a conditional expectation for Hilbert-valued random variables can be found in Bosq (2000). Under Model (1), it is easy to check that, for each $t$, $\text{E}(\epsilon_{t}|f_{t})={\bf 0}_{\mathcal{B}}$ is equivalent to $m(f)=\text{E}\left(f_{t+1}|f_{t}=f\right)$ (see Section B.1 in the Appendix). Our main challenge is to define for any given density $f$ an estimator $\widetilde{m}(f)$ of the unknown regression operator $m(f)=\text{E}\left(f_{t+1}|f_{t}=f\right)$ fulfilling density features (nonnegative functions that integrate to one) and good asymptotic behaviour while no reducing functional form (i.e. linearity, additivity, functional index model, etc) of the unknown regression operator $m(\cdot)$ is assumed.
@@ 2.2.2 Kernel estimator
With some smoothness properties, the functional form of $m(\cdot)$ is often estimated in a data-driven manner. There are a number of nonparametric functional estimators, such as the functional Nadaraya-Watson (NW) estimator (Ferraty & Vieu, 2006), the functional local linear estimator (Berlinet et al., 2011; Ferraty & Nagy, 2022), the functional $k$-nearest neighbour estimator (Burba et al., 2009, for the scalar response and Lian, 2011, for the functional response), the functional smoothing splines (Crainiceanu & Goldsmith, 2010), the functional wavelet estimator (Antoniadis & Sapatinas, 2003), the distance-based local linear estimator (Boj et al., 2010) and the functional neural network (Rossi & Villa, 2006). Throughout the paper, we introduce the functional Bayes NW estimator, i.e., the functional NW estimator in the Bayes-Hilbert space, because of its simplicity and mathematical elegance. The essence of functional Bayes NW kernel smoothing is to allow flexible estimation of the unknown regression operator. The functional Bayes NW estimator of the conditional mean can be defined as
@@
$\widetilde{m}_{N}(f)\,=\,\bigoplus^{N}_{t=1}w_{h_{\text{reg}}}(f_{t},\,f)\,\odot\,f_{t+1}\ \mbox{ with }\ w_{h_{\text{reg}}}(f_{t},\,f)\,=\,\frac{K_{\text{reg}}(h_{\text{reg}}^{-1}\,\|f_{t}\ominus f\|_{\mathcal{B}})}{\sum^{N}_{t=1}K_{\text{reg}}(h_{\text{reg}}^{-1}\,\|f_{t}\ominus f\|_{\mathcal{B}})},$ (2)
@@
where $K_{\text{reg}}(\cdot)$ is an asymmetric kernel function and $h_{\text{reg}}$ is a positive smoothing parameter (i.e. bandwidth). The subscript reg in the notation refers to the estimator of the regression operator. The vector space property of $\mathcal{B}_{2}(I)$ ensures that the Bayes NW estimator is a PDF. Depending on the kernel function and the bandwidth parameter, the nonparametric estimator in (2) is a weighted average of the observed densities and has two interesting properties: it fulfils the PDF constraints by construction, and due to its simplicity, its implementation is straightforward and computationally fast.
@@
In practice, for each $t=1,\ldots,N+1$, we observe the $n$ real random variable (rrv) $X_{t,1},\ldots,X_{t,n}$ that shares the same marginal density $f_{t}$. In other words, PDFs are not directly observable. This situation is particularly common in epidemiology. For example, suppose that at each date $t$, confirmed COVID-19 cases for all French departments are observed. In this case, $X_{t,j}$ is the COVID-19 confirmed cases at date $t$ of the department $j$, and the sample size $n$ is equal to the total number of departments. This is why our methodology first focusses on kernel density estimators $\widehat{f}_{1},\ldots,\widehat{f}_{N+1}$ of the $N+1$ probability densities $f_{1},\ldots,f_{N+1}$. In the second step, we just plug the estimated densities $\widehat{f}_{1},\ldots,\widehat{f}_{N+1}$ into the nonparametric density-on-density regression estimator (2), which provides our definitive kernel estimation:
$\widehat{m}_{N}(f)\,=\,\bigoplus^{N}_{t=1}w_{h_{\text{reg}}}(\widehat{f}_{t},\,f)\,\odot\,\widehat{f}_{t+1},$ (3)
@@
where $w_{h_{\text{reg}}}(\widehat{f}_{t},\,f)$ is defined in (2). A standard by-product of this estimating procedure in the time-series context is the following one-step-ahead sequential forecasting scheme. From the estimated density functions $\widehat{f}_{1},\ldots,\widehat{f}_{N+1}$, we compute $\widehat{\mkern-5.0mu\widehat{f}}_{N+2}=\widehat{m}_{N}(\widehat{f}_{N+1})$ which is the prediction of the density $f_{N+2}$. To forecast the next density, just consider the sample of random densities completed with the predicted one (i.e. $\widehat{f}_{1},\ldots,\widehat{f}_{N+1},\,\widehat{\mkern-5.0mu\widehat{f}}_{N+2}$) and build $\widehat{\mkern-5.0mu\widehat{f}}_{N+3}=\widehat{m}_{N+1}(\,\widehat{\mkern-5.0mu\widehat{f}}_{N+2})$. This one-step-ahead forecasting scheme can be iterated to obtain $\widehat{\mkern-5.0mu\widehat{f}}_{N+H}$ at any reasonable forecast horizon $H$.
2.2.3 Probabilistic setting and dependence structure